The magnetic field inside a 4.0-cm-diameter superconducting solenoid varies sinusoidally between 8.0 T and 12.0 T at a frequency of 6.0 Hz What is the maximum electric field strength at a point 1.0 cm from the solenoid axis? Express your answer to two significant figures and include the appropriate units. IA ?? ! XI) x-10n Units SubmitPrev Previous Answers Request Answer X Incorrect, Try Again; 3 attempts remaining Part B What is the value of B at the instant E reaches its maximum value? Express your answer to two significant figures and include the appropriate units. B-Value Units
Answer #1

v = \frac{\partial \phi }{\partial t}

v = \frac{\partial (B.A) }{\partial t}

v = A\frac{\partial (B) }{\partial t}

=v = \Pi R^{2} [ \frac{12-8}{\frac{1}{6}}]

R= 2cm= 0.02m

=0.030144 V

E= \frac{\partial v}{\partial x}

=\frac{0.030144}{0.01}

= 3 v/m